OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? a)3x - y = -11 b)3x + y = - 7 c)3x + y = 5 d)x + 3y = 3

OpenStudy (anonymous):

b

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = 2x - 1 and that pass through the point (4, 1). a)parallel: y = 2x - 7 perpendicular: y = -1/2 x + 3 b)parallel: y = 2x + 2 perpendicular: y = -1/2 x + 4 1/2 c)parallel: y = -1/2 x - 7 perpendicular: y = 2x + 3 d)parallel: y = -2x - 7 perpendicular: y =-1/2 x - 1

OpenStudy (anonymous):

A